# Unsupervised Prior Learning: Discovering Categorical Pose Priors from Videos

### v1

@@ 3.2 Connecting keypoints based on the connectivity prior
The connectivity of keypoints in objects is often fixed and rigid; for example, human arms maintain a relatively constant length, with a hand always connected to the torso via an arm and never connected to a foot. This rigidity in connectivity serves as a constraint, aiding in the regularization of pose estimation. In this section, we introduce the connectivity prior and explain how it can be used to regularize the connectivity strength between any pair of estimated keypoints in $T^{\prime}$ on $I$.
@@
Given the combined link map $S$ and the reference frame $I_{ref}$, PPL can reconstruct the image $I$. $I_{ref}$ shares the same background as $I$ but features a different foreground pose, thereby offering background information for reconstruction. Meanwhile, $S$ provides the connectivity details among all the estimated keypoints. This encourages PPL to learn accurate pose estimations while minimizing the influence of background noise. Therefore, we concatenate $S$ and $I_{ref}$ and feed them into a 2D-CNN to perform the image reconstruction $I_{recon}$, where $I_{recon}=\phi_{dec}(I_{ref},S)$.
@@ 3.3 Reconstruction of Keypoint Configuration with Memory
$M$ is a hierarchical memory storing compositional representations of prototypical poses. It contains $m$ memory banks $\{b_{1},b_{2},...,b_{m}\}$ of $k$ learnable tokens with $d$ randomly initialized parameters per token. $M$ is important to learn robust prior representations $V$ due to the following reasons: (1) By storing multiple prototypical poses, memory helps in aggregating these poses, which aids in creating a more robust and comprehensive prior that captures variations within an object category. (2) $M$ can assist in organizing information hierarchically, making it easier for PPL to retrieve relevant prototypical poses when making predictions in uncertain or ambugious scenarios, such as occlusion. and (3) By leveraging $M$, PPL can refine its predictions iteratively, using stored poses to adjust its outputs based on previously learned pose representations. Next, we introduce how $M$ is structured.
Given the estimated $N$ keypoints in $T^{\prime}$, we first encode it into $m$ tokens of dimension $d$ with several MLP-Mixer blocks (Tolstikhin et al., 2021), denoted as $MIX_{enc}$. We define $G$ as the collection of the $m$ tokens, where each token $g_{i}$ represents the embedding of a subset of the keypoints. Each token $g_{i}$ shares the same embedding space with the memory bank $b_{i}$ in memory $M$:
@@
If all the tokens in all the memory banks of $M$ learn to capture unique parts of prototypical poses, each $g_{i}$ should be able to retrieve the most similar token in the memory bank $b_{i}$ of $M$ and reconstructs $G$ itself. Here, we use $L2$ distance to measure the similarity between $g_{i}$ and $k$ token in $b_{i}$ of $M$. The token in $b_{i}$ that is most similar to $g_{i}$ is denoted as $g^{\prime}_{i}$. Thus, for each $g_{i}$, we can always find the corresponding $g^{\prime}_{i}$ from $b_{i}$, resulting in a collection of $G^{\prime}=[g^{\prime}_{1},g^{\prime}_{2},...,g^{\prime}_{m}]$. Another series of MLP-Mixer blocks (denoted as $MIX_{dec}$) is then used to decode $G^{\prime}$ back to $N$ keypoints $T^{\prime}_{recon}$, where $T^{\prime}_{recon}=MIX_{dec}(G^{\prime})$. See Figure LABEL:fig:mem_stru for the schematic.
Unlike PCT (Geng et al., 2023) where all tokens of a memory are within the same embedding space, We organize $M$ into $m$ memory banks, each containing $k$ tokens that represent an independent embedding space. This structure enables us to efficiently distill compositional parts from each memory bank, allowing us to form a general representation of a pose, which is distilled into the keypoint prior $T$. Specifically, we define $MP(\cdot)$ as the mean pooling operation. PPL pools one token vector among the $k$ vectors of each memory bank. These pooled $m$ vectors are further decoded by $MIX_{dec}$ into $N$ points of our distilled keypoint prior $T$. See Figure LABEL:fig:mem_cons for the schematic of distilling $T$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),...,MP(b_{m})])$.
@@ 3.4 Training and Inference
Our PPL is trained to jointly minimize all the four losses: the image reconstruction loss $L_{ir}$, the boundary loss $L_{b}$, the link regularization loss $L_{l}$, and the keypoint configuration reconstruction loss $L_{kr}$. We elaborate on these four losses below.
@@
$L_{b}=\sum_{*\in{x,y}}\begin{cases}|P^{\prime}_{i,*}|&\text{if }|P^{\prime}_{i,*}|>1,\\ 0&\text{otherwise}.\end{cases}\text{ and }L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1},$ (5)
@@
Link Regularization Loss. A person’s arm always maintains a fixed length regardless of the poses. Thus, we propose the constraint that links should be assigned a high weight if they do not vary significantly in length before and after the affine transformation. The loss $L_{l}$ encourages the preservation of link lengths during pose estimation. It is defined as in the equation above, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the affine transformation.
@@
Reconstruction Loss on Keypoint Configurations. In Section 3.3, given a collection of token representations in $G$, PPL retrieves the most similar tokens from each memory bank of the hierarchical memory $M$ and generates $G^{\prime}$ in a non-differentiable manner. To ensure that $M$ learns to store meaningful token embeddings that represent compositional parts of poses, the retrieved token representations $G^{\prime}$ from $M$ should closely match $G$. Moreover, if these compositional parts are structured correctly, the tokens should be able to decode into meaningful keypoint configurations $T^{\prime}_{recon}$ that are close to the original keypoint configurations $T^{\prime}$. Therefore, we introduce the keypoint configuration reconstruction loss $L{kr}$, defined as: $L_{kr}=\|T^{\prime}_{recon}-T^{\prime}\|_{2}+\|G-G^{\prime}\|_{2}$.
Training Techniques. To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE (Van Den Oord et al., 2017). Specifically, our PPL copies the gradients of $G^{\prime}$ to $G$ for backward propagation, allowing the gradients to flow through the quantization step. (2) The hierarchical memory $M$ is updated using an exponential moving average to smooth the gradient updates, particularly during the early stages of training when $G$ can be quite noisy. This approach helps stabilize the learning process and ensures that $M$ retains more reliable information over time. (3) For $M$ to learn effective representations of $G$, it requires an accurate estimation of $T^{\prime}$, which depends on a good prior $V$ distilled from $M$. This creates a chicken-and-egg problem that complicates training. To address this, we introduce two gradient detachments to separate the training processes. First, we detach the gradients from $T$ and train the keypoint transformation and image reconstruction pathway, as shown by the red arrows in Figure 2. Second, we detach the gradients from $T^{\prime}$ to train the memory encoder and decoder, $MIX_{\text{enc}}$ and $MIX_{\text{dec}}$, as indicated by the green arrows in Figure LABEL:fig:mem_stru.
Iterative Inference. We propose an iterative inference strategy (Figure 4). $4$ iterations are used for every experiment. In every iteration, we take the reconstructed frame $I_{recon}$ from the last iteration (the original frame $I$ for iteration 00) as the input. We infer its keypoints $T^{\prime}$ as the output keypoints of the current iteration. The hierarchical memory $M$ is used to reconstruct $T^{\prime}$ and the reconstructed keypoints $T^{\prime}_{recon}$ are used to obtain the reconstructed frame $I_{recon}$. $I_{recon}$ is then used as the input for the next iteration. We keep the original occluded frame $I$ as the reference frame for all the iterations. See Appendix A1 for more implementation details.
Figure 4: Overview of the iterative inference strategy in our PPL. During inference, we iteratively use the reconstructed frame $I_{recon}$ as input to estimate the pose $T^{\prime}$. The hierarchical memory $M$ refines the estimated pose $T^{\prime}$ and outputs $T^{\prime}_{recon}$. With the original test image $I$ as the reference frame $I_{ref}$, PPL reconstructs the image $I_{recon}$. It is then used as the input frame in the next iteration. See Section 3.4 for details.
4 Experiments
We use three video datasets: Human3.6m (Ionescu et al., 2013), Taichi (Siarohin et al., 2019), and YouTube dog videos with green backgrounds for our experiments. See Appendix A2 for details. On Human3.6m, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size. For Taichi, as in (He et al., 2022a; Siarohin et al., 2021; He et al., 2021; 2022b; Zhang et al., 2018) we use the summed $L2$ error computed at a resolution of $256\times 256$. To ensure a more accurate evaluation, unlike previous methods (He et al., 2022a; Schmidtke et al., 2021), we employ a minimum flow algorithm (Waissi, 1994) to align the predicted keypoints with the ground truth.
@@ 4 Experiments
Table 1: Keypoint detection on Human3.6m. We report the mean $L2$ error normalized by the image resolutions (Res.) of both $256\times 256$ and $128\times 128$. * is the $L2$ error calculated by our minimum flow algorithm. The best is in bold.
@@
Table 2: Keypoint detection on Taichi. For consistency with the baseline methods, we report the summed $L2$ error at a resolution of $256\times 256$. * is the $L2$ error calculated by our minimum flow algorithm. The best is in bold.
@@
4.1 Unsupervised Human Pose Estimation
@@
We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1 (Human3.6m) and Table 2 (Taichi). On both datasets, PPL significantly outperforms all baselines across all image resolutions. Among the baselines, AutoLink (He et al., 2022a) also incorporates learnable connectivity priors, similar to our PPL. However, its performance is inferior due to the absence of hierarchical memory. Additionally, we note that STT (Schmidtke et al., 2021), a baseline with human-defined priors, still underperforms compared to our PPL. Consistent with HPE (Yoo & Russakovsky, 2023), this suggests that pre-defined priors are not always optimal, and our PPL is able to learn more representative priors that outperform those manually defined.
@@
Visualization of Pose Estimation. We provide the visualization results of pose estimation on Human3.6m (Figure LABEL:fig:human-pose-vis) and YouTube dog videos (Figure LABEL:fig:dog-pose-vis). The visualization results demonstrate that PPL can learn priors and estimate poses of various object categories, without any external knowledge. For example, in Row 3, Column 2 of Figure LABEL:fig:human-pose-vis, PPL correctly estimates the bowing pose of a person. In Row 3, Column 5 of Figure LABEL:fig:dog-pose-vis, PPL correctly estimates the pose of a dog lowering down its head. Moreover, we found the quality of estimated poses of dogs is inferior to that of humans. This is primarily due to the significant morphological differences among various dog breeds. Additionally, dogs often perform actions such as turning around, which can lead to significant changes in pose that are difficult to accurately capture by our priors in 2D space.
@@
Visualization of the Pose Prior Changing with the Training Epochs. We visualize the progressively learnt pose priors by our PPL as a function of training epochs. Figure LABEL:fig:prior-vis illustrates that the keypoint prior converges to a human shape by the early stage of training (epoch 5). Notably, the learnable keypoints align with the human joints defined in the literature, and the connectivity among keypoints corresponds to the physical connections between body parts. As training continues, the connectivity prior gradually learns the skeletal structure of the human body, with irrelevant links between keypoints diminishing over time (as seen when comparing epochs 15 and 20).
@@ 4.2 Ablation Studies
Ablation on Number & Dimension of Tokens in Each Memory Bank. In our hierarchical memory, we used 34 memory banks. Here, we analyze the impact of the number of tokens per memory bank and the dimension of each token on PPL’s pose estimation performance. From Figure A1 in Appendix A3, we observed that PPL remains robust across different token counts and dimensions, although performance slightly improves with more tokens of higher dimensions. As a result, we fixed 16 tokens per memory bank, each with a dimension of 512, for all experiments.
Ablation on Number of Keypoints. We varied the number of keypoints in the pose priors from 4 to 32. The results in Figure A1 in Appendix A3 show that pose estimation accuracy improves as the number of keypoints in the prior increases. However, using 32 keypoints offers limited improvement compared to PPL with 16 keypoints for human pose estimation.
4.3 Pose Estimation in Occlusion Scene
To verify the robustness of PPL in occluded scenes, we divide the image into $32\times 32$ patches and apply two masking techniques: RandomMasking and CenterMasking. In RandomMasking, we randomly mask a certain proportion of image patches, with the proportion ranging from 0.1 to 0.4. In CenterMasking, we mask only the center region of the image, gradually increasing the masking size from $4\times 4$ to $12\times 12$ patches.
We explore the effect of occluded areas on PPL. From Figure 9 and Figure 9, we observe that at iteration 0, as the occluded areas increase, overall performance declines with larger occlusions. However, with our iterative inference strategy, PPL effectively infers the missing parts of the poses by utilizing prototypical poses stored in hierarchical memory and the learned priors. Notably, it restores partially occluded poses to reasonably complete full-body poses, leading to a lower $L2$ error, comparable to those without occlusion. This effect is more pronounced when the occlusion area is small or medium.
Visualization of Pose Estimation with Occlusion. We present the estimated poses by our PPL for occluded images as a function of the number of inference iterations in Figure LABEL:fig:occlusion-vis. Across both RandomMasking (top 2 rows) and CenterMasking (bottom 2 rows), with our iterative inference strategy, PPL successfully reconstructs the occluded part of the image after three iterations and meanwhile, predicts reasonable complete full-body poses.
Figure 8: PPL Results of Keypoint Detection as a Function of Number of Inference Iterations on Images with RandomMasking from Human3.6m. The “mask ratio” in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
Figure 9: PPL Results of Keypoint Detection as a Function of Number of Inference Iterations on Images with CenterMasking from Human3.6m. The “mask size” in the legend refers to the width and height of the masked region, on $1024$ image patches.

### v2

@@ 3.2 Connecting Keypoints Based on the Connectivity Prior
The connectivity of keypoints in objects is often fixed and rigid, for example, human arms maintain a relatively constant length, with a hand always connected to the torso via an arm and never connected to a foot. This rigidity in connectivity serves as a constraint, aiding in the regularization of pose estimation. In this section, we introduce the connectivity prior and explain how it can be used to regularize the connectivity strength between any pair of estimated keypoints in $T^{\prime}$ on $I$.
@@
Given the combined link map $S$ and the reference image $I_{ref}$, PPL can reconstruct the image $I$. $I_{ref}$ provides background information for reconstruction, while $S$ supplies foreground structural information by linking all the estimated keypoints with the connectivity prior. Therefore, we concatenate $S$ and $I_{ref}$ and feed them into a 2D-CNN to perform the image reconstruction $I_{recon}$, where $I_{recon}=\phi_{dec}(I_{ref},S)$.
@@ 3.3 Reconstruction of Keypoint Configuration with Memory
$M$ is a memory module storing compositional representations of prototypical poses. $M$ is hierarchical because it consists of $m$ memory banks $\{b_{1},b_{2},...,b_{m}\}$, each containing $k$ learnable vectors of dimension $d$. This setup allows for efficient retrieval and management of information at multiple levels, where $M$ can represent pose configurations while memory banks capture specific pose sub-structures. This hierarchy enables $M$ to effectively learn robust keypoint prior representations $T$ for the following reasons: (1) By storing multiple prototypical poses, memory helps in aggregating these poses, which aids in creating a more robust and comprehensive prior that captures variations within an object category. (2) $M$ can assist in organizing information hierarchically, making it easier for PPL to retrieve relevant prototypical poses when making predictions in uncertain or ambugious scenarios, such as occlusion. and (3) By leveraging $M$, PPL can refine its predictions iteratively, using stored poses to adjust its outputs based on previously learned pose representations. Next, we introduce how $M$ is structured.
Given the estimated $N$ keypoints in $T^{\prime}$, we first encode them into several tokens of dimension $d$ with several MLP-Mixer blocks (Tolstikhin et al., 2021), denoted as $MIX_{enc}$. The number of encoded tokens should always be equal to the number of memory banks, as we design these tokens to have different embedding space. Each token $g_{i}$ corresponds to the embedding of memory bank $b_{i}$ in memory $M$. We define $G$ as the collection of the $m$ tokens:
@@
If all the vectors in all the memory banks of $M$ learn to capture unique parts of prototypical poses, each $g_{i}$ should be able to retrieve the most similar vector in the memory bank $b_{i}$ of $M$ and reconstruct $G$ itself. Here, we use $L2$ distance to measure the similarity between $g_{i}$ and $k$ vectors in $b_{i}$ of $M$. The vector in $b_{i}$ that is most similar to $g_{i}$ is denoted as $g^{\prime}_{i}$. Thus, for each $g_{i}$, we can always find the corresponding $g^{\prime}_{i}$ from $b_{i}$, resulting in a collection of $G^{\prime}=[g^{\prime}_{1},g^{\prime}_{2},...,g^{\prime}_{m}]$. Another series of MLP-Mixer blocks (denoted as $MIX_{dec}$) is then used to decode $G^{\prime}$ back to $N$ keypoints $T^{\prime}_{recon}$, where $T^{\prime}_{recon}=MIX_{dec}(G^{\prime})$. See Figure ALABEL:fig:mem_stru in Appendix A5 for the schematic.
Unlike PCT (Geng et al., 2023) where all vectors of a memory are within the same embedding space, We organize $M$ into $m$ memory banks, each containing $k$ vectors that represent an independent embedding space. This structure enables us to efficiently distill each memory bank to form a general pose representation, which can be viewed as the keypoint prior $T$. Specifically, we define $MP(\cdot)$ as the mean pooling operation. PPL pools the $k$ vectors of each memory bank into one vector. These pooled $m$ vectors are further decoded by $MIX_{dec}$ into $N$ points of our keypoint prior $T$. See Figure ALABEL:fig:mem_cons in Appendix A5 for the schematic of distilling $T$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),...,MP(b_{m})])$.
Compared to simply using a moving average of the learned poses, the advantages of employing hierarchical memory and keypoint configuration reconstruction are as follows: (1) the hierarchical memory increases the complexity of the embedding space with compositional parts, enhances the expressive capability for poses and facilitates more accurate pose retrieval for occluded images; (2) the compositional nature of the hierarchical memory addresses data bias, enabling the distillation of a less biased prior.
@@ 3.4 Training and Inference
Our PPL is trained to jointly minimize all the four losses: the image reconstruction loss $L_{ir}$, the boundary loss $L_{b}$, the link regularization loss $L_{l}$, and the keypoint configuration reconstruction loss $L_{kr}$. We elaborate on these four losses below. Additionally, we use three training techniques for stable convergence of the network. We present the details of these training techniques in Appendix A6.
@@
$L_{b}=\sum_{*\in{x,y}}\begin{cases}|P^{\prime}_{i,*}|&\text{if }|P^{\prime}_{i,*}|>1,\\ 0&\text{otherwise}.\end{cases}$ (5)
@@
Link Regularization Loss. A person’s arm always maintains a fixed length regardless of the poses. Thus, we propose the constraint that links should be assigned a high weight if they do not vary significantly in length before and after the affine transformation. The loss $L_{l}$ encourages the preservation of the link lengths during pose estimation. It is defined as in the equation below, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the affine transformation:
$L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1}.$ (6)
@@
Reconstruction Loss on Keypoint Configurations. In Section 3.3, given a collection of token representations in $G$, PPL retrieves the most similar vectors from each memory bank of the hierarchical memory $M$ and generates $G^{\prime}$ in a non-differentiable manner. To ensure that $M$ learns to store meaningful token embeddings that represent compositional parts of poses, the retrieved vectors $G^{\prime}$ from $M$ should closely match $G$. Moreover, if these compositional parts are structured correctly, the vectors should be able to decode into meaningful keypoint configurations $T^{\prime}_{recon}$ that are close to the original keypoint configurations $T^{\prime}$. Therefore, we introduce the keypoint configuration reconstruction loss defined as: $L_{kr}=\|T^{\prime}_{recon}-T^{\prime}\|_{2}+\|G-G^{\prime}\|_{2}$.
@@
Table 1: Keypoint detection on CUB-200-2011. We report the mean $L2$ error normalized by the image resolutions of $128\times 128$. We use 10 keypoints for CUB-Aligned and 4 keypoints for CUB-001, CUB-002, CUB-003, and CUB-all. The best is in bold.
Method CUB- aligned CUB- 001 CUB- 002 CUB- 003 CUB- all
(Zhang et al., 2018) 5.36 26.9 27.6 27.1 22.4
(He et al., 2021) 5.21 22.6 29.1 21.2 14.7
(He et al., 2022b) 3.23 22.1 22.3 21.5 12.1
AutoLink (He et al., 2022a) 3.51 20.2 19.2 18.5 11.3
PPL (ours) 3.19 19.3 18.6 17.3 10.5
Table 2: Keypoint detection on Human3.6m. We report the mean $L2$ error normalized by the image resolutions (Res.) of both $256\times 256$ and $128\times 128$. * is the result trained by using masked images as reference images. The best is in bold.
@@
Table 3: Keypoint detection on Taichi. For consistency with the baseline methods, we report the summed $L2$ error at a resolution of $256\times 256$. * is the result trained by using masked images as reference images. The best is in bold.
@@
Iterative Inference. We propose an iterative inference strategy (Figure 3). $4$ iterations are used for every experiment. In every iteration, we take the reconstructed image $I_{recon}$ from the last iteration (the original image $I$ for iteration 00) as the input. We infer its keypoints $T^{\prime}$ as the output keypoints of the current iteration. The hierarchical memory $M$ is used to reconstruct $T^{\prime}$ and the reconstructed keypoints $T^{\prime}_{recon}$ are used to obtain the reconstructed image $I_{recon}$. $I_{recon}$ is then used as the input for the next iteration. We keep the original occluded image $I$ as the reference image for all the iterations. See Appendix A1 for more implementation details.
Figure 3: Overview of the iterative inference strategy in our PPL (Section 3.4). During inference, we iteratively use the reconstructed image $I_{recon}$ as input to estimate the pose $T^{\prime}$. The hierarchical memory $M$ refines the estimated pose $T^{\prime}$ and outputs $T^{\prime}_{recon}$. With the original image $I$ as the reference image $I_{ref}$, PPL reconstructs the image $I_{recon}$. It is then used as the input image in the next iteration.
4 Experiments
For quantitative experiments, we use three image datasets: Human3.6m (Ionescu et al., 2013), Taichi (Siarohin et al., 2019b, a), and CUB-200-2011 (Wah et al., 2011). For fair comparisons, we set the number of keypoints $N$ to be the same as (He et al., 2022a). For qualitative visualizations, we use Youtube dog videos, Flowers (Nilsback & Zisserman, 2008), 11k-Hands (Afifi, 2019), and Horses (Zhu et al., 2017). See Appendix A2 for details. We utilize video frames as $I_{ref}$ for the Human3.6m, Taichi, and YouTube dog videos, while randomly masked images are used as $I_{ref}$ for other datasets. On Human3.6m and CUB-200-2011, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size. For Taichi (He et al., 2022a; Siarohin et al., 2021; He et al., 2021, 2022b; Zhang et al., 2018), we use the summed $L2$ error computed at a resolution of $256\times 256$.
4.1 Unsupervised Categorical Pose Estimation
@@
We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1 (CUB-200-2011), Table 2 (Human3.6m), and Table 3 (Taichi). On all datasets, PPL significantly outperforms all baselines across all image resolutions. Among the baselines, AutoLink (He et al., 2022a) also incorporates learnable connectivity priors. However, its performance is inferior due to the absence of hierarchical memory. Notably, even when masked images are used as $I_{ref}$, PPL consistently outperforms AutoLink. This demonstrates that the effectiveness of PPL is primarily attributed to its pose prior, rather than relying on cross-frame reconstruction. Additionally, we note that STT (Schmidtke et al., 2021), a baseline with human-defined priors, still underperforms PPL. Consistent with HPE (Yoo & Russakovsky, 2023), this suggests that pre-defined priors are not always optimal. PPL can learn more representative priors that outperform those manually defined.
@@
Visualization of Pose Estimation. We provide the visualization results of pose estimation on Human3.6m (Figure ALABEL:human-vis), YouTube dog videos (Figure ALABEL:dog-vis), 11k-Hands (Figure ALABEL:hand-vis), CUB-200-2011 (Figure ALABEL:bird-vis), Horse (Figure ALABEL:horse-vis), and Flowers (Figure ALABEL:flower-vis) in Appendix A7. The visualization results demonstrate that PPL can learn categorical priors and estimate poses for various object categories, without any external annotations. For example, in Row 2, Column 2 of Figure ALABEL:human-vis, PPL correctly estimates the bowing pose of a person. In Row 2, Column 5 of Figure ALABEL:dog-vis, PPL correctly estimates the pose of a dog lowering down its head. Moreover, we found the quality of estimated poses of dogs is inferior to that of humans. This is primarily due to the significant morphological differences among various dog breeds. Additionally, dogs often perform actions such as turning around, which can lead to significant changes in pose that are difficult to accurately capture by our priors in 2D space. Additional visualization results for Flowers, Hands, and Horses in Appendix A7 further demonstrate that our PPL can be applied for various categories. Notably, flowers are rigid objects with no degrees of freedom like animals. Yet, our PPL still learns meaningful categorical priors for them.
@@
Visualization of the Pose Prior Changing with the Training Epochs. We visualize the progressively learnt pose priors by our PPL as a function of training epochs. Figure LABEL:fig:prior-vis illustrates that the keypoint prior converges to a human shape by the early stage of training (epoch 5). Notably, the learnable keypoints align with the human joints defined in the literature, and the connectivity among keypoints corresponds to the physical connections between body parts. As training continues, the connectivity prior gradually learns the skeletal structure of the human body, with irrelevant links between keypoints diminishing over time, as seen when comparing epochs 15 and 20.
@@ 4.2 Ablation Studies
Ablation on Number & Dimension of Vectors in Each Memory Bank. In our hierarchical memory, we used 34 memory banks. Here, we analyze the impact of the number of vectors per memory bank and the dimension of each vector on PPL’s pose estimation performance. From Figure 5 in Appendix A3, we observed that PPL remains robust across different vector counts and dimensions, although performance slightly improves with more vectors of higher dimensions. As a result, we fixed 16 vectors per memory bank, each with a dimension of 512, for all experiments.
Ablation on Number of Keypoints. We varied the number of keypoints in the pose priors from 4 to 32. The results in Figure 5 in Appendix A3 show that pose estimation accuracy improves as the number of keypoints in the prior increases. However, using 32 keypoints offers limited improvement compared to PPL with 16 keypoints.
4.3 Pose Estimation in Occluded Scenes
To verify the robustness of PPL in occluded scenes, we divide the image into $32\times 32$ patches and apply two masking techniques: RandomMasking and CenterMasking. In RandomMasking, we randomly mask a certain proportion of image patches, with the proportion ranging from 0.1 to 0.4. In CenterMasking, we mask only the center region of the image, gradually increasing the masking size from $4\times 4$ to $12\times 12$ patches. We explore the effect of occluded areas on PPL. From Figure A2 and Figure A2 in Appendix A4, we observe that at iteration 0, as the occluded areas increase, overall performance declines with larger occlusions. However, with our iterative inference strategy, PPL effectively infers the missing parts of the poses by utilizing prototypical poses stored in hierarchical memory and the learned priors. Notably, it restores partially occluded poses to reasonably complete full-body poses, leading to a lower $L2$ error, comparable to those without occlusion. This effect is more pronounced with smaller occluded areas.
Visualization of Pose Estimation with Occlusion. We present the estimated poses by our PPL for occluded images as a function of the number of inference iterations in Figure LABEL:fig:occlusion-vis. Across both RandomMasking and CenterMasking, with our iterative inference strategy, PPL successfully reconstructs the occluded image parts after three iterations and meanwhile, predicts reasonable full-body poses.